A 425N force is applied horizontally to a 33 kg box on a flat table, if the acceleration of the box is 5 m/s2 what is the coefficient of friction between the box and the table?
1
Expert's answer
2020-11-17T13:02:46-0500

Let's apply the Newton's Second Law of Motion:


"F_{appl}-F_{fr}=ma,""F_{appl}-\mu mg=ma,""\mu=\dfrac{F_{appl}-ma}{mg},""\mu=\dfrac{425\ N-33\ kg\cdot 5\ \dfrac{m}{s^2}}{33\ kg\cdot 9.8\ \dfrac{m}{s^2}}=0.80"

Answer:

"\mu=0.80"